OpenStudy (anonymous):

The largest population the forest can sustain is represented mathematically by the limit as t approaches infinity. Determine this limit. P(t)= 20/ 1+3e^(-0.02t) where p is squirrel population in hundreds, and t is time in weeks

OpenStudy (anonymous):

so, I definitely hope this will be true- well , I am just a student:\[\lim_{t \rightarrow \infty}\frac{ 20 }{ 1+ 3e ^{-0.02t} }\] =\[\lim_{t \rightarrow \infty}\frac{ 20(1-3e ^{-0.02t}) }{ (1+3e ^{-0.02t})(1-3e ^{-0.02t}) }\] =\[\lim_{t \rightarrow \infty}\frac{ 20-60e ^{-0.02t} }{ 1-9e ^{-0.04t} }\] = \[\lim_{t \rightarrow \infty}\frac{ 20- 0 }{ 1-0 } =20\] this is so because \[3e ^{-0.02t }= \frac{ 3 }{e ^{-0.02t} }\] and as t gets to infinity, the whole expression becomes zero. The same goes for the denominator

OpenStudy (anonymous):

Yup thats right thanks for helping :)

OpenStudy (anonymous):

hey, am so happy!